# Physics and Deep Learning in Computational Wave Imaging

### v1

@@ IV-A2 Semi-supervised Methods
Fully supervised learning is a straightforward approach to applying ML to CWI problems, but it is often difficult to obtain an adequate number of labeled samples. The acquisition of unlabeled waveform data, in contrast, is usually less challenging. For instance, in seismic FWI, distributed acoustic sensing (DAS) [46] enables efficient collection of a large volume of unlabeled acoustic data. Various semi-supervised models have been developed [47, 48, 49, 50] to alleviate the scarcity of labeled samples.
Specifically, Gómez et al. [47] illustrate how semi-supervised learning can be adapted to solving CWI. The high-level idea is to adapt the training set to the particular waveform $\bm{d}$ that was measured, by incrementally generating more data samples that are close to the measured data. The first step is to learn a reconstruction operator $g_{0}$ from the waveform-data manifold to velocity-map manifold using the original pairwise data. From the measured data, use $g_{0}(\bm{d}+\eta)$ to create a new batch of models $\bm{m}$, and then apply $f$ to those models to create a new batch of labeled samples. The training set is augmented with the new labeled samples, and a new regression $g_{1}$ is fit. With each iteration, more labeled samples are in the regime of the measured sample, and the performance of the resulting $g$ is improved for the waveform $\bm{d}$ of interest.
In general, semi-supervised methods can be viewed as a strategy for generating “pseudo labels” using unlabeled data or other auxiliary information. For example, the above work directly derives velocity maps from unlabeled waveform data to support a supervised learning approach. Similarly, Feng et al. [49] utilize available well log profiles—which are detailed recordings of geological formations encountered during drilling—to generate additional “pseudo labels” for velocity maps. In contrast, Cai et al. [51] also produce “pseudo labels” from unlabeled waveform data, but they employ a Cycle-GAN architecture to facilitate the learning process. Sun et al. [50] generate “pseudo labels” by extrapolating low-frequency velocity components, which serve as initial estimates for conventional physics-based CWI approach.
@@ IV-A3 Self-Supervised Methods
Self-supervised learning (SSL), which exploits both labeled and unlabeled data, provides an alternative to alleviate the requirements of a large volume of high-quality labels. SSL is a machine-learning paradigm leveraging both labeled and unlabeled data. Its key idea is to allow models to learn from measurements themselves without the need for a vast amount of labels. An approach, named “Auto-Linear”, developed by Feng et al. [10] use SSL to separately train an encoder and decoder, each in their domains, and further observes a near-linear relationship across domains in the latent spaces. The network structure of Auto-Linear is illustrated in Fig. 3b. It consists of two pairs of encoder and decoder: one for waveform data denoted as $\mathcal{E}_{d}$ and $\mathcal{D}_{d}$, respectively; and the other one for velocity maps denoted as $\mathcal{E}_{m}$ and $\mathcal{D}_{m}$. Mathematically, Auto-Linear can be posed as
$\widehat{\bm{m}}$ $=g(\bm{d})=\mathcal{D}_{m\theta^{*}_{dm}}(u_{\theta^{*}}(\mathcal{E}_{d\theta^{*}_{ed}}(\bm{d})))\;,$ (5)
$\mathrm{s.t.}\;\theta^{*}$ $=$
$\operatorname*{arg\,min}_{\theta}$ $\sum_{\bm{m},\bm{d}\in\Phi}\mathcal{L}(\mathcal{D}_{m\theta^{*}_{dm}}(u_{\theta}(\mathcal{E}_{d\theta^{*}_{ed}}(\bm{d})))),\bm{m}),$ (6)
$\{\theta^{*}_{em}(\Phi),\theta^{*}_{dm}(\Phi)\}$ $=$
$\operatorname*{arg\,min}_{\theta_{em},\theta_{dm}}$ $\sum_{\bm{m},\_\in\Phi}\mathcal{L}(\mathcal{D}_{m\theta_{dm}}(\mathcal{E}_{m\theta_{em}}(\bm{m})),\bm{m}),$ (7)
$\{\theta^{*}_{ed}(\Phi),\theta^{*}_{dd}(\Phi)\}$ $=$
$\operatorname*{arg\,min}_{\theta_{ed},\theta_{dd}}$$\sum_{\_,\bm{d}\in\Phi}\mathcal{L}(\mathcal{D}_{d\theta_{dd}}(\mathcal{E}_{d\theta_{ed}}(\bm{d})),\bm{d}),$ (8)
where $\theta$ represents the trainable parameters of the corresponding network components, $\mathcal{L}(\cdot,\cdot)$ is a loss function, and $\Phi$ is the training dataset with paired samples of $(\bm{m},\bm{d})$. As indicated by the loss functions above, Auto-Linear is trained in a two-step process: self-supervised learning of the separate encoder-decoder pairs (i.e., Eqs. (IV-A3,IV-A3)) followed by fully supervised learning of the relationship between their latent spaces (i.e., Eq. (IV-A3)).
SSL is an emerging area within CWI, still in an early stage of development. This approach has been successfully applied to extract useful features from waveform data and velocity maps, as demonstrated in studies by Gao et al. [52], Liu et al. [53], and Feng et al. [10]. Research has shown that SSL can be seamlessly integrated with the forward modeling operator as in Eq. (2). Notably, Liu et al. [53] develop a physics-driven, self-supervised method for CWI that uses mostly unlabeled data supplemented by a small set of labeled data. This technique significantly reduces the need for extensive labeled data, a common pitfall of fully supervised methods, and improves generalization by incorporating physical principles. Conversely, Gao et al. [52] explore SSL within a traditional physics-based CWI framework, designing a network that identifies and extrapolates frequency envelope features across different frequency bands. These features are then utilized to compute the physics misfit function, as in Eq. (3). Their approach proves effective in capturing low-frequency data components, thus enhancing CWI methods in overcoming the challenges associated with ill-posedness.
IV-A4 Unsupervised Methods
In the most extreme case, no labels are available. For instance, in USCT, acquiring ultrasound waveforms is generally straightforward and does not require prior knowledge of the target velocity maps. Similarly, in seismic FWI, DAS offers an efficient means of data collection. To overcome such a challenge, there are methods built on a completely unsupervised learning strategy to solve various CWI problems [54, 11, 55]. Here, we will describe the method Unsupervised Physics-informed Full-Waveform Inversion (UPFWI), developed by Jin et al. [11], as an example to demonstrate how to leverage physics knowledge to alleviate the requirement for high-quality label information. As illustrated in Fig. 3c, the UPFWI model consists of two components: an encoder-decoder network and a full-wave equation. Mathematically, UPFWI can be posed as a minimization problem
@@ IV-A4 Unsupervised Methods
$\mathrm{s.t.}\;\bm{\theta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\sum_{\bm{d}_{i}\in\bm{\Phi}_{u}}\mathcal{L}(f(g_{\bm{\theta}}(\bm{d}_{i})),\,\bm{d}_{i})\,,$
where $\bm{\Phi}_{u}$ represents an unsupervised dataset that contains waveform data $\bm{d}_{i}$ alone. Similarly to its supervised counterpart shown in Fig. 3a, UPFWI takes waveform measurements as input and produces velocity maps as output. The major difference is that, since labels are not available, UPFWI defines the loss function in the data space rather than in the velocity space. To do so, the velocity model estimated by the encoder-decoder network is mapped to measurement data by propagating the wave-equation forward model. Numerically, the continuous full-wave equation is discretized using finite differences and incorporated in the loss function as a differentiable operator. The rationale for this is that, similar to model-based FWI, a good match between the measured and simulated waveform indicates a reasonable reconstruction of the physical parameters from the encoder-decoder network. Another perspective is to consider the full-wave equation as an additional layer of regularization [11].
Broadly, unsupervised learning strategies can be divided into two categories: per domain and per sample methods. Per domain methods offer the advantages of fast “on-line” reconstruction (i.e. use a pretrained network) but may suffer from bias induced by the training samples. per sample methods do not require any training sample at all; however, incur “on-line” reconstruction costs comparable to conventional model-based CWI. Specifically, the UPFWI method, as described by Eq. (9), is categorized under per domain because its network weights $\bm{\theta}^{*}$ are optimized for a specific subdomain $\bm{\Phi}_{u}$. Works such as those by Sun et al. [55], Zhu et al. [56], and He et al. [57] exemplify per sample methods, which rely on parameterization techniques [58]. Further details on per sample methods will be covered in the section below titled Parameterization Solution Methods.
Figure 4: The schematic illustration of the network architecture of (a) Fourier-DeepONet, a simulations/measurements-driven method [59], (b) invertible network, a physics-aware method [60], and (c) parameterization-based method [56].
@@ IV-B Machine Learning Methods by Learning Strategy
CWI, unlike typical ML problems such as natural language processing, is governed by underlying physics. We are thus presented with an opportunity to incorporate critical physics and prior knowledge to potentially improve the performance of ML models in solving CWI problems [3]. As illustrated in Fig. 2, different approaches have been developed to incorporate physics into ML; we distinguish the following four:
@@ IV-B1 Simulations/Measurements-Driven Methods
Full-physics simulations and measurements from experiments represent a major source of data for many ML methods for solving physics problems including CWI. Here, simulation data refers to synthetic waveforms generated by virtual imaging. This is achieved by applying the forward model in Eq. (1) to various realizations of the medium properties spanning the physical space of interest. The measurements are these data being acquired through lab or controlled experimental setup. Particularly, given the significant recent increase in available computing power, a very large number of pairwise realizations and waveforms becomes available, enabling fully supervised models as provided in Eq. (4). Here, we use the model (called “Fourier-DeepONet”) developed by Zhu et al. [59] as an example to illustrate how simulations can be utilized to train ML models for solving CWI problems as shown in Fig. 4a. Fourier-DeepONet was developed to enhance the generalization ability of waveform sources including both frequency range and source locations. It builds on two major components: DeepONet (deep operator networks) [61] and U-FNO (a block combining Fourier neural operator and U-net) [62]. Particularly, source parameters (i.e., frequencies and source locations) are employed as the inputs of the trunk net to facilitate the data-driven model with varying source frequencies and locations. The branch net and trunk net encode the waveform data and source parameters, respectively. They are linear transformations lifting inputs to high dimensional space. A few U-FNO blocks were then utilized as decoders to produce high-resolution imaging output.
@@
Fully supervised learning strategies, as discussed in the section on Fully Supervised Methods, prove highly effective in utilizing the large volume of simulation-data pairs. Beyond the above-mentioned method, a substantial body of research has explored simulation/measurement driven methods for FWI [1, 41, 42, 63, 64], USCT [65, 9, 66, 67, 68], and NDT [69, 70, 43]. However, a common issue with these approaches is their limited generalization ability; that is, the models’ suboptimal predictability when faced with unseen data significantly different from those in the training set. Thus, additional effort (such as transfer learning, active learning, etc.) [71, 72, 47] may be required to overcome the distribution shift between simulations and real-world measurements. Another method that has proven effective in enhancing generalization is the “CNN Correction after Direct Inversion” approach, as proposed by Jin et al. [73]. This method initially applies model-based inversion to closely approximate the physics model of the system, producing an estimate near the true solution. Subsequently, a CNN utilizes this estimate along with high-quality labels to learn how to eliminate artifacts and thus improve reconstruction quality. Jeong et al. [68] implement a similar strategy to address the USCT problem.

### v2

@@ IV-A2 Semi-supervised Methods
Figure 4: Schematic illustration of semi-supervised learning strategies in CWI. In the CycleGAN-based low-frequency extrapolation method [56], the network learns to translate between unpaired high-frequency field data and synthetic low-frequency data using adversarial and cycle-consistency losses. Image courtesy of the authors of [56].
Fully supervised learning provides a direct means to apply machine learning to CWI, but its utility is limited by the scarcity of labeled data. Acquiring velocity models as labels can be costly and time-consuming, whereas collecting unlabeled waveform data is typically much more feasible. For example, in seismic FWI, distributed acoustic sensing (DAS) [57] allows efficient acquisition of large volumes of unlabeled seismic data. To address the imbalance between labeled and unlabeled data availability, various semi-supervised learning approaches have been developed [58, 59, 60, 56].
Among these, [56] proposes a CycleGAN-based SSL framework designed to extrapolate missing low-frequency content in field seismic data. The method is trained using unpaired data from two domains: band-limited field shot gathers (typically 4–10 Hz), denoted $h\in\mathcal{H}$, and synthetic low-frequency data (0–4 Hz), denoted $l\in\mathcal{L}$, derived from known subsurface models. Rather than relying on paired training data, the model learns mappings between domain distributions using adversarial learning and cycle-consistency constraints. The architecture comprises two generators, $G_{HL}:\mathcal{H}\rightarrow\mathcal{L}$ and $G_{LH}:\mathcal{L}\rightarrow\mathcal{H}$, along with two discriminators, $D_{L}$ and $D_{H}$, trained jointly with adversarial, identity, and cycle-consistency losses. Waveform structure is preserved by enforcing the cycle-consistency conditions:
$G_{LH}(G_{HL}(h))\approx h\quad\text{and}\quad G_{HL}(G_{LH}(l))\approx l,$
allowing the network to produce plausible low-frequency extrapolations even in the absence of ground-truth labels. Once trained, $G_{HL}$ is applied to real field data to generate low-frequency components, which are then used to enhance inversion quality in a subsequent two-stage full-waveform inversion (FWI) pipeline. The method is successfully validated on both synthetic and real marine field datasets.
An alternative strategy in semi-supervised learning is to generate “pseudo-labels” from unlabeled data or auxiliary measurements. For instance, [58] uses physics-based models to infer velocity maps from raw waveform data, effectively creating supervision signals for a downstream learning task. Similarly, [60] utilizes available well-log data to construct pseudo-labels for velocity profiles. [61] also generates pseudo-labels from waveform data but incorporates a CycleGAN architecture to improve the fidelity of the inferred labels. Together, these methods highlight the flexibility of SSL in balancing data availability and model generalization in CWI.
@@ IV-A3 Self-Supervised Methods
Figure 5: Schematic illustration of the network architecture of SiameseFWI [62], a self-supervised learning framework for full waveform inversion. The model uses a Siamese neural network to compare simulated and observed seismic data in a shared latent space, enhancing sensitivity to key structural features. Image courtesy of the authors of [62].
Self-supervised learning (SSL) involves the model generating its own supervision signals from unlabeled data. The model creates tasks, such as predicting parts of the data from other parts, where labels are implicitly generated through the data itself. Its key idea is to allow models to learn from measurements themselves without the need for a vast amount of labels. As an example, SiameseFWI [62] has been proposed as a novel self-supervised framework for FWI that uses a Siamese neural network to enhance the comparison between simulated and observed seismic data. Unlike traditional pixel-wise loss functions such as the $\ell_{2}$ norm, SiameseFWI transforms both data types into a shared latent space via two identical convolutional neural networks with shared weights. This design enables the network to focus on meaningful features (e.g., P-wave arrivals), and it minimizes the Euclidean distance between the learned representations to guide velocity model updates. The method achieves robust inversion performance on synthetic benchmarks and field data from Western Australia, demonstrating greater accuracy and resilience to noise, poor initial models, and unknown source wavelets.
SSL remains an emerging area within CWI, with growing interest in its application to low-frequency extrapolation, which is an essential step in overcoming the lack of low-frequency content in seismic data that limits FWI performance [14]. One early effort [63] introduces an SSL method that trains a neural network using paired seismic datasets with shared high-frequency content but varying low-frequency components. The model learns to reconstruct missing low-frequency information, although its performance is inherently constrained by the frequency range present in the original data. Furthermore, the required downsampling can degrade high-frequency resolution, limiting time precision. To address these challenges, [64] proposes a more effective two-stage SSL framework for low-frequency extrapolation. The warm-up stage initializes training using high-pass filtered inputs and original data as pseudo-labels. In the subsequent iterative data refinement stage, the model refines its predictions by using outputs from the previous epoch as pseudo-labels, with their high-pass filtered versions serving as new inputs. This iterative refinement progressively closes the gap between predictions and the ideal low-frequency content, enhancing extrapolation accuracy and robustness.
IV-A4 Unsupervised/Unlabeled Methods
Figure 6: Schematic overview of two representative unsupervised/unlabeled methods. (a) A CycleGAN-based unsupervised framework for low-frequency reconstruction in FWI, which learns to map high-frequency inputs to low-frequency components without paired labels [65]. (b) The UPFWI architecture, a physics-informed unlabeled method that maps waveform data to velocity models by minimizing the mismatch between observed and simulated waveforms via a differentiable wave equation [11]. Image (a) courtesy of the authors of [65].
This category encompasses methods that operate without paired supervision; i.e., where only waveform samples are available and no velocity labels are required. It includes both unsupervised learning techniques that extract structure or features from data distributions, and unlabeled inversion methods that reconstruct medium properties by enforcing data consistency through physical models.
Unsupervised learning methods typically involve training neural networks on tasks defined from the data itself, allowing them to uncover latent structure or reconstruct missing information. For example, [65] proposes an unsupervised learning framework for low-frequency reconstruction in FWI using cycle-consistent adversarial networks (CycleGAN) as shown in Fig. 6a. Their method aims to infer large-scale low-frequency content from high-frequency seismic inputs, recovering information often absent in real field data. Unlike supervised methods that require labeled training pairs, CycleGAN enables training directly on unpaired seismic data, which is more practical for real-world applications. To address training instability and enhance output quality, the authors constructed synthetic targets combining high-, mid-, and low-frequency components, and incorporated additional loss functions to improve reconstruction performance. This design allows the model to supply critical low-frequency components for more robust downstream inversion.
In contrast, unlabeled methods focus on mapping waveform data to velocity models by directly minimizing the mismatch between observed and simulated waveforms, using the full wave equation as a physical constraint. A representative example is the Unsupervised Physics-informed Full-Waveform Inversion (UPFWI) framework (shown in Fig. 6b) developed in [11]. UPFWI couples a neural encoder-decoder architecture with a differentiable wave-equation simulator, enabling end-to-end training with a loss defined in the data space:
@@ IV-A4 Unsupervised/Unlabeled Methods
$\mathrm{where}\;\bm{\theta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\sum_{\bm{d}_{i}\in\bm{\Phi}_{u}}\mathcal{L}(f(g_{\bm{\theta}}(\bm{d}_{i})),\,\bm{d}_{i})\,.$
Here, $\bm{\Phi}_{u}$ denotes a dataset of unlabeled waveform measurements, $g{\bm{}\theta}$ is the network mapping waveforms to velocity models, and $f$ represents the forward modeling operator. The simulated waveforms are compared to the observed data to guide the network’s learning. Methods in this category can be further divided into two subtypes:
  • Per-sample methods, such as those in [66] and [67], reparameterize the model and optimize its parameters individually for each input, mimicking traditional inversion routines.
  • Per-domain methods, like UPFWI [11] and [68], train a single model across a dataset and apply it to similar examples, enabling faster inference but possibly inheriting dataset-specific biases.
Together, these unsupervised and unlabeled strategies reflect a growing effort to decouple learning-based inversion from the need for explicit supervision, while still leveraging physical constraints and inductive priors to achieve accurate reconstructions.
IV-A5 Comparative Overview of Supervising Strategies for CWI
Fully Supervised Learning.  Fully supervised learning provides the most direct application of machine learning to CWI by training models on datasets composed of paired waveform measurements and corresponding physical property maps. This strategy has been explored across various domains—including geophysics, medical imaging, and materials characterization—to approximate complex mappings from time-domain signals to spatial distributions of physical parameters such as velocity, impedance, or elastic moduli. Its appeal lies in its conceptual simplicity and fast inference: once trained, these models can generate real-time predictions without iterative solvers or numerical simulations.
However, the theoretical appeal of this approach belies practical challenges. The need to implicitly learn the underlying physics from data alone makes these models highly data-intensive. They often require orders of magnitude more labeled examples than physics-informed or hybrid approaches to achieve comparable generalization. Label acquisition is especially difficult in CWI contexts, where high-fidelity ground-truth models are either unavailable (as in seismic exploration), invasive to obtain (as in biomedical applications), or destructive (as in nondestructive testing of materials). Moreover, supervised models are vulnerable to domain shift: models trained on synthetic or idealized data may not transfer well to field conditions unless the training data closely mimics real-world complexity.
Despite these limitations, there are notable cases where fully supervised learning has achieved success, typically under carefully designed conditions. For example, [6] demonstrates accurate ultrasonic defect reconstruction in multilayered materials by tailoring the synthetic training data to match the sensor setup and propagation environment used in experiments. Similarly, [69] trains a Transformer-based network on over 20 million examples from a curated dataset of velocity models, enabling their model to generalize across diverse geophysical structures. These successes underscore that fully supervised approaches can work when domain fidelity and data scale are carefully managed—but they also highlight the inherent difficulty of deploying such methods directly in practical scenarios.
Ultimately, while fully supervised learning serves as a foundational paradigm for data-driven CWI, its dependency on exhaustive labeled datasets and lack of embedded physics limits its standalone applicability. These limitations have led to increased interest in semi-supervised, self-supervised, and physics-informed frameworks that can leverage unlabeled data or incorporate domain knowledge to improve generalization and interpretability.
Semi-Supervised Learning (SSL). Semi-supervised learning represents a pragmatic approach to bridging the gap between the data efficiency of supervised learning and the flexibility of unlabeled methods. It is particularly suited for CWI applications, where waveform measurements are abundant but reliable ground-truth velocity models are difficult to obtain. The key strength of SSL lies in its ability to exploit the information embedded in unlabeled data to guide or regularize model training, thereby enhancing generalization with limited supervision. Unlike fully supervised methods, SSL frameworks can be designed to integrate physics-based constraints or exploit structural relationships within the data to generate pseudo-labels or consistent feature mappings. In CWI, this enables networks to extract meaningful representations even in the absence of dense annotations. When properly configured, SSL methods can achieve performance close to their fully supervised counterparts, while being more robust to overfitting and more scalable to large, real-world datasets. By accommodating both labeled and unlabeled sources, SSL provides a flexible and data-efficient strategy for learning in complex wave-based imaging tasks.
Self-Supervised Learning. Self-supervised learning is an emerging paradigm that enables models to learn meaningful representations directly from unlabeled data by formulating surrogate tasks in which the labels are derived from the data itself. In the context of CWI, self-supervised learning offers a powerful means to exploit the inherent structure and redundancy in waveform measurements, without relying on explicit supervision. Recent advances have demonstrated its potential in tasks such as low-frequency extrapolation and feature embedding. For example, the SiameseFWI framework [62] introduces a contrastive learning setup where pairs of simulated and observed waveforms are mapped into a shared latent space using a Siamese neural network, enabling robust misfit evaluation without the need for pixel-wise loss functions. Other approaches leverage signal masking or predictive reconstruction tasks to derive supervisory signals, allowing the network to learn frequency-domain or spatial-temporal patterns directly from the input data. The core advantage of self-supervised methods lies in their ability to harness large-scale unlabeled datasets for pretraining, which can subsequently enhance performance on downstream inversion tasks, particularly when labeled data is scarce. However, the success of self-supervised learning depends heavily on the design of pretext tasks and alignment with the physics of wave propagation, making it an active and promising area for further exploration in CWI.
Unsupervised/Unlabeled Methods. Unsupervised/unlabeled learning strategies provide a label-efficient alternative for tackling CWI problems by either extracting latent patterns directly from data or leveraging physics-based constraints to supervise the learning process. In unsupervised approaches, the goal is to uncover intrinsic structures within waveform data—such as temporal or frequency-domain features—that can facilitate inversion, without requiring paired velocity models. These methods are particularly effective when used for tasks like representation learning, data clustering, or generative modeling, and can serve as a powerful pretraining step for downstream inversion. In contrast, unlabeled physics-informed methods directly embed physical models, such as the wave equation, into the learning objective, allowing end-to-end inversion from measurements to physical parameters without explicit supervision. These techniques are especially advantageous in applications where unlabeled waveform data are abundant, but acquiring ground-truth models remains difficult. Despite the benefit of bypassing labeled data, these methods often require careful regularization, reliable forward modeling, and principled training strategies to ensure convergence and physical plausibility. Together, unsupervised and unlabeled methods offer a flexible and scalable paradigm for model training under realistic data constraints.
@@ IV-B Machine Learning Methods by Learning Strategy
Computational wave imaging, unlike typical ML problems such as natural language processing, is governed by underlying physics. We are thus presented with an opportunity to incorporate critical physics and prior knowledge to potentially improve the performance of ML models in solving CWI problems [3]. As illustrated in Fig. 2, different approaches have been developed to incorporate physics into ML; we distinguish the following four:
@@ IV-B1 Simulations/Measurements-Driven Methods
Figure 7: Schematic illustration of Fourier-DeepONet, a simulations/measurements-driven method [70]. Image courtesy of the authors of [70].
@@
Full-physics simulations and measurements from experiments represent a major source of data for many ML methods for solving physics problems including CWI. Here, simulation data refers to synthetic waveforms generated by virtual imaging. This is achieved by applying the forward model in Eq. (1) to various realizations of the medium properties spanning the physical space of interest. The measurements are these data being acquired through lab or controlled experimental setup. Particularly, given the significant recent increase in available computing power, a very large number of pairwise realizations and waveforms becomes available, enabling fully supervised models as provided in Eq. (5). These simulation/measurement-driven methods learn the underlying physics indirectly through data, by observing how changes in physical parameters affect the resulting wavefields. In this way, they do not encode physics through explicit equations, but rather through exposure to physics-rich data generated via forward modeling.
Here, we use the model (called “Fourier-DeepONet”) developed by [70] as an example to illustrate how simulations can be utilized to train ML models for solving CWI problems as shown in Fig. 7. Fourier-DeepONet is a representative example of this class because it explicitly leverages frequency-domain information embedded in simulation data to drive learning. It was developed to enhance the generalization ability of waveform sources including both frequency range and source locations. It builds on two major components: DeepONet (deep operator networks) [71] and U-FNO (a block combining Fourier neural operator and U-net) [72]. Particularly, source parameters (i.e., frequencies and source locations) are employed as the inputs of the trunk net to facilitate the data-driven model with varying source frequencies and locations. The branch net and trunk net encode the waveform data and source parameters, respectively. They are linear transformations lifting inputs to high dimensional space. A few U-FNO blocks were then utilized as decoders to produce high-resolution imaging output. This use of frequency-domain learning makes Fourier-DeepONet particularly well-suited to handle variability in waveform sources, thus exemplifying how simulation-driven approaches can exploit structured physics knowledge implicitly embedded in data.
@@
Fully supervised learning strategies, as discussed in the section on Fully Supervised Methods, prove highly effective in utilizing the large volume of simulation-data pairs. Beyond the above-mentioned method, a substantial body of research has explored simulation/measurement driven methods for FWI [1, 51, 52, 73, 74], USCT [75, 9, 76, 77, 78], and NDT [79, 80, 53]. However, a common issue with these approaches is their limited generalization ability; that is, the models’ suboptimal predictability when faced with unseen data significantly different from those in the training set. Thus, additional effort (such as transfer learning, active learning, etc.) [81, 82, 58] may be required to overcome the distribution shift between simulations and real-world measurements. Another method that has proven effective in enhancing generalization is the “CNN Correction after Direct Inversion” approach, as proposed by [83]. This method initially applies model-based inversion to closely approximate the physics model of the system, producing an estimate near the true solution. Subsequently, a CNN utilizes this estimate along with high-quality labels to learn how to eliminate artifacts and thus improve reconstruction quality. [78] implement a similar strategy to address the USCT problem.